OpenStudy (anonymous):

solving linear equations 6(y+2)-4=10

OpenStudy (anonymous):

do you know how to distribute the 6 through the parenthesis ?

OpenStudy (anonymous):

hello ? Are you there ?

OpenStudy (anonymous):

let me know when you want to participate and I will help you

OpenStudy (anonymous):

i think i wanna participate now :)

OpenStudy (anonymous):

lol...ok do you know how to distribute the 6 through the parenthesis ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do that and let me see what you get :)

OpenStudy (anonymous):

do you need help doing that ?

OpenStudy (anonymous):

does this help : 6*y + 6*2 - 4 = 10

OpenStudy (anonymous):

the answer is -3

OpenStudy (anonymous):

6(y + 2) - 4 = 10 6y + 12 - 4 = 10 6y + 8 = 10 6y = 10 - 8 6y = 2 y = 1/3 check... 6(1/3 + 2) - 4 = 10 6(2/6 + 12/6) - 4 = 10 6(14/6) - 4 = 10 14 - 4 = 10 10 = 10 (correct) y = 1/3

OpenStudy (anonymous):

do you understand ?